# Bayesian Multivariable Bidirectional Mendelian Randomization

### v1

@@ 5 Conclusion
In this work, We introduce MR.RGM, a Bayesian multivariable, bidirectional Mendelian randomization framework that performs network-wide causal inference. The method explicitly accommodates unmeasured confounding and feedback (cyclic) structure, and it jointly selects and estimates (i) the directed causal network among traits, (ii) instrument-trait effects, and (iii) a residual covariance whose off-diagonal elements indicate latent confounding. Sparsity-inducing priors yield interpretable graphs and effect maps, while the fully Bayesian formulation provides uncertainty quantification for edges, effects, and confounding links. The inference can be carried out from sufficient statistics, so the method works seamlessly with summary-level data when individual-level records are unavailable.
@@
Extensive simulations with realistic network structure show that MR.RGM consistently outperforms competing approaches, delivering higher graph-recovery AUC and lower effect-estimation error as sample size grows. In horizontal pleiotropic regimes, MR.RGM+ achieves the best performance for both graph and effect recovery and attains near-perfect accuracy in SNP-gene selection at moderate to large sample size. The confounding structure recovered from the estimated residual covariance attains AUCs close to one in large samples, demonstrating that the same posterior output simultaneously supports causal network estimation and latent-confounder mapping. Despite using MCMC, the framework is computationally competitive, scaling to problem sizes typical in gene regulatory network analyses (i.e., sizes of typical signaling pathways) within practical runtimes.
Applications to GTEx v7 skeletal muscle tissue and the OneK1K B-cell group further validate the approach: we recover high-probability causal edges concordant with known biology and reveal plausible latent confounding modules, all with principled uncertainty measures. Taken together, these results position MR.RGM as a powerful computational tool for multivariable MR in modern transcriptomic studies.
Supplementary Material
@@
R package availability: The full implementation of our Bayesian MR framework is publicly available as the MR.RGM package, which includes functions for posterior inference and graph visualization. The package can be accessed via:

### v2

@@ 5 Conclusion
We have introduced MR.RGM, a Bayesian multivariable, bidirectional Mendelian randomization framework that performs network-wide causal inference and structure learning. The method explicitly accommodates unmeasured confounding and feedback (cyclic) structure, and it jointly selects and estimates (i) the directed causal network among traits, (ii) instrument-trait effects, and (iii) a residual covariance whose off-diagonal elements indicate latent confounding. Sparsity-inducing priors yield interpretable graphs and effect maps, while the fully Bayesian formulation provides uncertainty quantification for edges, effects, confounding links, and higher-order network motifs. The inference can be carried out from sufficient statistics, so the method works seamlessly with summary-level data when individual-level records are unavailable.
@@
Extensive simulations with realistic network structure show that MR.RGM consistently outperforms competing approaches, delivering higher graph-recovery AUC and lower effect-estimation error as sample size grows. In horizontal pleiotropic regimes, MR.RGM+ achieves the best performance for both graph and effect recovery and attains near-perfect accuracy in SNP-gene selection at moderate to large sample size. The confounding structure recovered from the estimated residual covariance attains AUCs close to one in large samples, demonstrating that the same posterior output simultaneously supports causal network estimation and latent-confounder mapping. Despite using MCMC, the framework is computationally competitive, scaling to sizes of typical signaling pathways within a practical runtime.
Applications to GTEx v77 skeletal muscle tissue and the OneK11K B-cell group further validate the approach: we recover high-probability causal edges including feedback loops concordant with known biology and reveal plausible latent confounding modules, all with principled uncertainty measures. Beyond individual edges, the proposed NetworkMotif function enables posterior inference on higher-order regulatory structures, such as feedback loops, feedforward loops, and cascades, yielding interpretable motif-level probabilities that summarize coordinated signaling mechanisms supported by existing biological literature and quantify uncertainty.
There are some remaining practical considerations. When individual-level measurements of $(\mathbf{X,Y,U})$ are available, MR.RGM can be applied directly; if covariates $\mathbf{U}$ are unavailable, the algorithm can be used without them, as implemented in our software. For summary-level analyses, the method requires the matrices $(\mathbf{S_{yy},S_{yx},S_{xx},S_{yu},S_{xu},S_{uu}})$. In practice, covariate-related quantities such as $\mathbf{S_{yu}}$, $\mathbf{S_{xu}}$, and $\mathbf{S_{uu}}$ are sometimes unavailable; in such cases, the method can be run without covariates, as in the individual-level setting. The trait–trait covariance $\mathbf{S_{yy}}$ can be reliably estimated as long as each pair of traits is measured in at least one study, even if different pairs come from different cohorts. A common limitation is that $\mathbf{S_{yx}}$ may exclude trans-SNP associations because many GWAS only report genome-wide significant hits; this is inherent to summary-level MR and implies that the method performs best when a reasonably dense set of SNP–trait associations is available. When in-sample LD ($\mathbf{S_{xx}}$) is unavailable and external reference LD must be used, population mismatch may introduce LD misspecification. This can be mitigated by selecting a reference panel matched on ancestry, or by modeling $\mathbf{S_{xx}}$ as a latent positive-definite matrix—e.g., with a Wishart or inverse-Wishart prior centered at the observed reference LD, so that LD uncertainty is propagated into the posterior. Finally, although the present work focuses on continuous traits (e.g., gene expression), the framework can be extended to binary or categorical complex traits using standard data-augmentation strategies such as Pólya–Gamma augmentation for logistic models or probit-based latent-variable formulations.
Taken together, these strengths and considerations position MR.RGM as a flexible and powerful computational tool for network-wide Mendelian randomization in modern genomic and transcriptomic studies, supporting inference at both the edge and motif levels, with clear avenues for extension to complex traits and large-scale biobank analyses.
Data and Code Availability
@@
R package availability. The full implementation of the proposed Bayesian Mendelian randomization framework is publicly available as the MR.RGM R package. The package includes functionality for posterior inference, causal graph estimation, uncertainty quantification, and network motif analysis, and can be accessed via:
Reproducibility of simulations and real data analyses. All simulation code and analysis scripts used to generate the results in this paper are available at:
The repository contains scripts to reproduce all simulation studies and real data analyses, along with detailed documentation and software version information.